OpenStudy (anonymous):

The sum of two numbers is 67. The larger number is 4 more than 2 times the smaller. What are the two numbers?

OpenStudy (anonymous):

x = larger number, y = smaller numer x + y =67 2y + 4 = x If you substitute you can solve for both x and y

OpenStudy (anonymous):

2y + 4 + y = 67 3y + 4 = 67 minus 4 on both sides 3y = 63 divide by 3 on both sides y = 21 Substitute "y" back in the original equation (x + y =67) x + 21 = 67 minus 21 on both sides x = 46

OpenStudy (anonymous):

46 is 4 more than 2 times the smaller number 21, and 46 + 21 = 67. I hope this helped

OpenStudy (anonymous):

it did thx @ZDoubleM